OpenStudy (anonymous):

Exponential decay: a farmer in china discovers a mammal hide that contains 30% of its original c-14. find the age of the hide to the nearest year. N=N(sub o)e^-kt N(subo)= initial amount of C-14 (at time=o) N=amount of C-14 at time t t=time in years k=0.0001

OpenStudy (anonymous):

you are given everything you need to set up the equation right?

OpenStudy (anonymous):

N=N(sub o)e^-kt N(subo)= initial amount of C-14 (at time=o) N=amount of C-14 at time t t=time in years k=0.0001

OpenStudy (anonymous):

\(N=30\%=.3, k=.0001\) so your job is to solve \[.3=e^{-.000t}\] for \(t\) which takes the following two steps. write in equivalent logarithmic form as \[\ln(.3)=-.000t\] and then divide both sides by \(-.0001\) to get \[t=\frac{\ln(.3)}{-.0001}\] and the final step is to use a calculator to see what the heck this number is

OpenStudy (anonymous):

thank yu.

OpenStudy (anonymous):

yw